### v1

# sEMG-Driven Physics-Informed Gated Recurrent Networks for Modeling Upper Limb Multi-Joint Movement Dynamics

Abstract
Exoskeletons and rehabilitation systems offer great potential for enhancing human strength and recovery through advanced human-machine interfaces (HMIs) that adapt to movement dynamics. However, the real-time application of physics-informed neural networks (PINNs) is limited by their reliance on fixed input lengths and surrogate models. This study introduces a novel physics-informed Gated Recurrent Network (PiGRN) designed to predict multi-joint torques using surface electromyography (sEMG) data. The PiGRN model employs a Gated Recurrent Unit (GRU) to convert time-series sEMG inputs into multi-joint kinematics and external loads, which are then integrated into an equation of motion to ensure consistency with physical laws. Experimental validation with sEMG data from five participants performing elbow flexion-extension tasks showed that the PiGRN model accurately predicted joint torques for 10 unfamiliar movements, with RMSE values between 4.02% and 11.40% and correlation coefficients ranging from 0.87 to 0.98. These findings highlight the PiGRN’s potential for real-time exoskeleton and rehabilitation applications. Future research will explore more diverse datasets, improve musculoskeletal models, and investigate unsupervised learning methods.
Keywords Surface electromyography  $\cdot$ Joint dynamics  $\cdot$ Human machine interface $\cdot$ Gated Recurrent Unit, Physics-informed neural network  $\cdot$ Inverse dynamics
1 Introduction
In recent decades, wearable robotic assistive devices have significantly evolved, boosting human power and aiding rehabilitation [1]. Upper extremity exoskeletons, for instance, enhance strength and endurance, enabling users to exceed their usual capabilities, making them promising for assisting personnel in heavy item transport and aiding elderly and disabled individuals in rehabilitation [2, 3]. A vital aspect of these devices is the human-machine interface (HMI), which aims to understand human behavior for effective control [4], by accurately detecting human movement dynamics. This allows devices to respond naturally to the user’s movements [5]. Modeling human movement dynamics involves parameters such as joint angles, angular velocities, angular accelerations, joint forces, and torques. While direct sensing and tracking of these parameters using sensors can be effective, it may introduce delays and limited flexibility for assistive devices [6, 7], as these methods detect movements only after they occur, failing to provide real-time assistance.
Surface electromyography (sEMG) shows promise as a natural HMI, measuring neuromuscular electrical potentials noninvasively and predicting movement dynamics beforehand [8]. Physics-based neuromusculoskeletal (NMS) models have used sEMG to develop effective exoskeleton control strategies based on joint torques [9, 10, 11]. However, challenges such as complex muscle geometry modeling and the need for intricate calibration limit its real-time applicability in assistive device control.
Researchers are using artificial intelligence to link sEMG signals with joint movement dynamics, employing regression-based machine learning. Neural networks, like dynamic RNNs and ANNs, have shown effectiveness in analyzing muscle activity and limb movement, predicting various parameters from sEMG signals, such as muscle force, limb kinematics, and kinetics [12, 13, 14, 15, 16, 17]. Research has further refined these methods for better sEMG-to-joint angle mapping and developing virtual human models [18, 19, 20]. Recent advancements also include the exploration of individual deep learning models such as CNNs, RNNs, autoencoders, and LSTMs, as well as their combinations, for estimating limb movement dynamics from multi-channel EMG signals [21, 22, 23, 24]. Training complex deep learning models can be a time-consuming process, particularly when dealing with large datasets, leading to substantial costs. Additionally, pre-trained models frequently do not perform well when applied to new subjects’ data [25]. Transfer learning is suggested for tasks like predicting muscle forces and estimating limb kinematics for new subjects with limited data [26, 27]. However, these deep learning methods tend to act as ’black boxes,’ often ignoring the underlying physics of neuromusculoskeletal processes.
Advancements in musculoskeletal (MSK) modeling and joint kinematics employ physics-informed machine learning techniques, enhancing flexibility and computational efficiency while adhering to physical principles [28, 29]. One study integrated MSK modeling into a CNN-based physics-informed model to map EMG signals to muscle forces and joint kinematics [30]. Zhang et al. further personalized the MSK model using CNN-based transfer learning and reduced training time through a distributed framework [27, 31]. Taneja et al. used an ANN-based physics-informed parameter identification neural network to predict motion and identify parameters in MSK systems from raw sEMG signals and joint motion data [32]. Another study combined MSK physics principles with a GAN to estimate muscle force and joint kinematics from sEMG signals [33]. Ma et al. proposed an ANN-based physics-informed neural network approach to predict muscle forces and identify muscle-tendon parameters using unlabeled sEMG data [34]. Despite these advancements, few challenges remain in integrating physics and machine learning in MSK modeling.
The above mentioned studies utilized deep learning models based on ANN or CNN, which have a limitation: they require fixed input sizes, posing challenges for real-time applications [30, 27, 31, 34]. sEMG signals, being time-series data with variable lengths, conflict with this requirement, making ANN or CNN models unsuitable for dynamic input data structures. Muscle force prediction was their key focus, often determined using surrogate models like the CMC tool from OpenSim [35]. While useful for computing ground-truth values, surrogate models have drawbacks. They may not provide precise label data, leading to erroneous predictions, and their modeling and simulation can be costly and time-consuming, along with challenges in data recording. Additionally, most studies focused on single-joint movements, neglecting the transferability of models trained on sEMG data across different muscle loading scenarios. Understanding how variations in EMG signal strength affect model performance is crucial for diverse loading scenarios.
In recent years, the Gated Recurrent Unit (GRU), introduced by Kyunghyun Cho et al. in 2014 [36], offers a simplified alternative to traditional RNNs, resembling LSTM networks [37]. GRUs streamline structure reduces parameter count while maintaining performance. They are proficient in analyzing time series data, GRUs excel in capturing long dependencies without vanishing gradient issues [38, 39]. Furthermore, their recursive design enables them to sustain and update a hidden state across time, efficiently handling sequential data and capturing temporal patterns [40]. Hence GRUs are ideal for processing sEMG signal time series, representing muscle activity patterns effectively.
This work aims to achieve three objectives:
  1. I
    Analysis of sEMG for Predicting Joint Torque: We propose a novel physics-informed Gated Recurrent Network (PiGRN) to analyze sEMG signals to predict joint torques. Also, the GRU within the PiGRN extracts temporal features from sEMG time series, accommodating variable input sequences.
  2. II
    Correlating sEMG with Joint Kinematics and External Load: In the proposed framework the GRU component first maps sEMG signals to joint kinematic parameters (angle, velocity, acceleration) and external load, while adhering with dynamic motion equations. This integration serves as a soft constraint on predicted kinematics and load.
  3. III
    Validation Across Multiple Degrees of Freedom and Load: We validated the robustness of PiGRN model in predicting joint torques across multiple degrees of freedom, considering varied muscle loading scenarios in an upper limb model. This includes accurately predicting shoulder and elbow joint torques during elbow FE tasks using sEMG recordings from specified muscles under different loads.
The rest of this article proceeds as follows: Section II details the methodology, including GRU architecture, inverse dynamics equations, PiGRN architecture, and combined loss function design. Section III explains Systems Validation, covering experimental setup, data processing, feeding, and training, along with PiGRN model hyperparameters, torque prediction, and evaluation criteria. Section IV presents results, while Section V discusses future work, and Section VI concludes.
2 Methodology
2.1 The Gated Recurrent Unit (GRU) Architecture
The Figure 1 illustrates the architecture of a Gated Recurrent Unit (GRU), comprising two gates: an update gate ($z_{t}$) and a reset gate ($r_{t}$). These gates play a crucial role in regulating the flow of information within the network. Within this framework, as depicted in Figure. 2, the recursive nature enables the model to effectively retain and update information across sequential sEMG data. At each time step, the GRU receives an input sEMG vector $x_{t}$ and combines it with the previous hidden state $h_{t-1}$ to compute an intermediate state. The update gate determines how much of the previous memory to retain by computing $z_{t}$ using $h_{t-1}$, the previous hidden state, and $x_{t}$ the current input. The reset gate decides how much of the past information to forget using $r_{t}$. The candidate hidden state $h^{\prime}_{t}$ is then computed using $r_{t}$, $h_{t-1}$, and $x_{t}$. Finally, the new hidden state $h_{t}$ combines the previous hidden state $h_{t-1}$ and the candidate hidden state $h^{\prime}_{t}$ to predict joint kinematic parameters and external load sequence. Here, the key equations that govern the information flow within a GRU is explained.
Figure 1: The GRU framework
$Update\hskip 2.84544ptGate:z_{t}=\sigma\*(W_{z}*x_{t}+U_{z}*h_{t-1})$
$Reset\hskip 2.84544ptGate:r_{t}=\sigma\*(W_{r}*x_{t}+U_{r}*h_{t-1})$
$Candidate\hskip 2.84544ptHidden\hskip 2.84544ptState:h^{\prime}_{t}=\tanh\*(W*x_{t}+U*r_{t}*h_{t-1})$
$Hidden\hskip 2.84544ptstate:h_{t}=z_{t}*h_{t-1}+(1-z_{t})*h^{\prime}_{t}$
$Output:\hat{y_{t}}=W^{y}*h_{t}$
Here, $W_{i}$ and $U_{i}$ are the weight matrices corresponding to the current input $x_{t}$ and the previous hidden state $h_{t-1}$ respectively, with $i$ being $z$ and $r$. The GRU architecture utilizes the sigmoid function, $\sigma$, and the hyperbolic tangent function, $\tanh$, as activation functions. Additionally, $W$ is another weight matrix used in the computation of the candidate hidden state.
Figure 2: The recursive model of GRU at time instance $t$
2.2 Inverse Dynamics Equation of Motion
This work utilizes the inverse dynamics equation of motion method to estimate joint torques in the musculoskeletal (MSK) systems. Specifically, we employ a previously derived equation of motion for elbow flexion-extension (FE) motion in the sagittal plane, considering an external load ($m$) held in the hand. The FE motion involves two degrees of freedom (DOF) at the shoulder and elbow joints, modeled as revolute joints. The MSK comprises the upper arm (humerus) and the lower arm (ulna and radius), treated as a rigid segments. Body segment mass and center of mass are estimated following methods described in Herman et al. [41], while segment moment of inertia (MOI) is calculated using regression equations from Hinrichs et al. [42]. The equation of motion is represented by Equation 1.
$I(q(t),m){\ddot{q}(t)}+{C((q(t),m)},{\dot{q}(t)})+{{G(q(t),m})}={\tau(t)}$ (1)
where $q(t)$, ${\dot{q}}(t)$, ${\ddot{q}}(t)$ are the joint kinematics parameters including angle, angular velocities, and angular accelerations, respectively; $I(q(t),m)$ is the system inertia matrix; $C((q(t),m),{\dot{q(t)}})$ is the centrifugal and Coriolis force matrix; $G(q(t),m)$ gravitational force on the system all as a function of external load ($m$).
2.3 Main Framework of Physics Informed Gated Recurrent Network (PiGRN)
Figure 3 illustrates the main framework of the proposed method for training the Physics Informed Gated Recurrent Network (PiGRN) using varied load sEMG signals. The GRU takes sEMG signals as input and predicts joint kinematic parameters, including joint angles ($q(t)$), angular velocities (${\dot{q}}(t)$), and angular accelerations (${\ddot{q}}(t)$) for all joints and external load ($m$) in the musculoskeletal (MSK) model.
Figure 3: Physics Informed Gated Recurrent Network Architecture (PiGRN), here, $N_{T}$ denotes length of time series data, $t$ denotes time instance. $N_{C}$ and $N_{dof}$ denotes number of channels of sEMG signal and total number of degrees of freedom in the MSK system , $h_{i-1}$ and $GRU_{t}$ denotes hidden state and GRU block at at time instance $t$. $\lambda_{p}hysics$ is physics loss weighting factor which is a hyperparameter. Losses are based on MSE i.e. Mean square error.
Here a GRU model is employed to extract temporal dependencies in muscle activation patterns and predict joint kinematics from time-series EMG data. Initially, sEMG signals recorded under varied loads are input into the GRU, which then predicts joint kinematics. These predictions are then used in the MSK equation to calculate predicted torque, compared with actual torque to determine physics loss. This physics loss is incorporated into the GRU’s loss function, ensuring a combined loss that minimizes both data loss and physics-based loss. Ultimately, this personalized network effectively maps sEMG signals to joint kinematics.
2.4 Design of Combined Loss Function
To train the proposed PiGRN, the framework utilizes a combined loss function comprising two main components: the standard data loss (representing supervised learning) and the physics-based loss (representing residual loss). The data loss, calculated as Mean Squared Error (MSE) loss, aims to minimize the discrepancy between the ground truth and the predicted values. On the other hand, the physics-based loss ensures adherence to physical constraints governed by the equation of motion, achieving this by minimizing the residual of joint torque. The components of the loss functions in PiGRN are as follows: The data loss (supervised learning loss):
$L_{\text{data}}=L_{q}+L_{\dot{q}}+L_{\ddot{q}}+L_{m}$ (2)
where, the total data loss, denoted as $L_{\text{data}}$, is obtained by summing $L_{q}$, $L_{\dot{q}}$, $L_{\ddot{q}}$, and $L_{m}$, which represents the losses for joint angle, angular velocity, angular acceleration, and external load respectively. Each of these data loss functions is defined as the mean squared error between the predicted and target values, as given in Equation 3.
$L_{\text{u}}=\frac{1}{N}\sum_{i=1}^{N}\left|(\hat{u}(x_{i});{\hat{\theta}})-u_{i}\right|^{2}$ (3)
where $\mathcal{D}=\{(x_{i},u_{i})\}_{i=1}^{N}$ is a set of measured data points with $x_{i}$ as sEMG inputs and $u_{i}$ as corresponding joint kinematics values, and $\hat{u}(x_{i})$ is the network’s prediction at $x_{i}$ with trainable parameters $\hat{\theta}$. The physics loss (Residual Loss):
$L_{\text{physics}}=\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}((\hat{q}(x_{i}),\hat{\dot{q}}(x_{i}),\hat{\ddot{q}}(x_{i}),\hat{m}(x_{i}));{\hat{\theta}})\right|^{2}$ (4)
where $\mathcal{F}$ represents the residual of the joint torque, $x_{i}$ are sEMG data points, and $N$ is the number of data points. The combined loss function: The total loss function $L_{\text{total}}$ in a PiGRN is a weighted sum of these individual loss components:
$L_{\text{total}}=\lambda_{\text{data}}L_{\text{data}}+\lambda_{\text{physics}}L_{\text{physics}}$ (5)
where $\lambda_{\text{data}}$, and $\lambda_{\text{physics}}$ are hyperparameters controlling the relative importance of each term, here the value of $\lambda_{\text{data}}$ is one.
The neural network is trained by minimizing the total loss function $L_{\text{total}}$, Equation 5. This involves adjusting the network parameters $\hat{\theta}$ so that the predicted solution $(\hat{q}(x_{i});{\hat{\theta}}),(\hat{\dot{q}}(x_{i});{\hat{\theta}}),(\hat{\ddot{q}}(x_{i}));{\hat{\theta}})$, and $(\hat{m}(x_{i});{\hat{\theta}})$ fits the observed data, while also adhering to the physical law represented by the equation of motion of the MSK system dynamics during movements as described in Equation 6.
$\begin{split}\mathcal{F}(\hat{q}_{i},\hat{\dot{q}}_{i},\hat{\ddot{q}}_{i},\hat{m}_{i};{\hat{\theta}})=I(\hat{q}_{i},\hat{m}_{i};\hat{\theta})\hat{\ddot{q}}_{i}+C((\hat{q}_{i},\hat{m}_{i};\hat{\theta}),\hat{\dot{q}}_{i})\ldots\\ \ldots+G(\hat{q}_{i},\hat{m}_{i};\hat{\theta})-\tau\end{split}$ (6)
Here, $\mathcal{F}(\hat{q}_{i},\hat{\dot{q}}_{i},\hat{\ddot{q}}_{i},\hat{m}_{i};{\hat{\theta}})$, represents residuals related to Equation 6 for the $i^{th}$ sample. $(\hat{q}(x_{i});\hat{\theta}$), $(\hat{\dot{q}}(x_{i});\hat{\theta}$), $(\hat{\ddot{q}}(x_{i});\hat{\theta}$), and $(\hat{m}(x_{i});\hat{\theta}$) refers to the joint kinematics and the external load predicted by the GRU model with trainable parameters $\hat{\theta}$ for the $i^{th}$ sample. The most accurate predictions are achieved by minimizing the combined loss function $J$, as defined in Equation 7, for the optimal GRU parameters denoted as ${\theta}$.
${\theta}=\underset{\hat{\theta}}{\operatorname{argmin}}\,(J),\\ \quad\text{where}\;J=L_{\text{data}}+\lambda_{\text{physics}}L_{\text{physics}}$ (7)
3 SYSTEM VALIDATION PROCEDURES
3.1 Experimental setup, Data collection and Data processing
3.1.1 Experimental setup
The study enlisted five healthy male volunteers, aged $26\pm 3$ years, with an average body mass of $73\pm 4$ kg and a height of $174\pm 3.5$ cm.
Figure 4: Joint torque estimation flow chart once the PiGRN model is trained
These participants engaged in elbow flexion-extension (FE) exercises utilizing 0 kg, 2 kg, and 4 kg weights, enabling movement along the sagittal plane for both the elbow and shoulder joints. Concurrent measurements of surface electromyography (sEMG) signals and joint angle data were recorded. Approval for the experimental procedure was obtained from the All India Institute of Medical Sciences (AIIMS) in New Delhi. The sEMG data were captured using Noraxon’s Ultium EMG sensor system (4000 Hz), while the elbow joint angle was tracked employing the Noraxon Ninox camera system (125 Hz).
3.1.2 Data collecion
The study recorded sEMG signals and joint angles from five participants. Four electrodes were placed on each participant’s biceps long head, biceps short head, triceps long head, and triceps lateral head muscles, following guidelines from [43]. Additionally, four reflective markers tracked shoulder and elbow joint angles. Each trial began with a screen cross and beep, signaling motion initiation, and ended manually. A two-second rest followed each elbow flexion-extension (FE). Participants completed 12 runs, with 4 runs each for 0 kg, 2 kg, and 4 kg loads, comprising 10 FE trials per run.
3.1.3 Data pre-processing
We enhanced joint angle data by applying a Gaussian filter with a 10-standard deviation and a 6-window size. For sEMG data, we first filtered signals between 10-450 Hz to remove noise, then converted negative amplitudes to positive through full-wave rectification. Next, a 7 Hz low-pass filter smoothed the signal, yielding linear envelopes. These were normalized by dividing them by the maximum voluntary contraction (MVC) EMG values from recorded trials, ensuring consistency across trials and subjects. Given EMG signals’ 4000 Hz frequency and joint angle recordings’ 125 Hz, we downsampled the EMG envelope by 32 to match joint angle data points.
3.2 Data feeding and Training of PiGRN
The dataset was divided into a 3:1 ratio for training and testing. Angular velocity and acceleration were computed from joint angles using the central difference method. Data was collected in sets of 10 elbow flexion-extension (FE) movements, with rest between each repetition. To speed up training and simplify data handling, we segmented the data into complete FE movement repetitions, each comprising 800 timesteps, including rest. Without rest, each FE repetition was approximately 450 timesteps long. The GRU network takes sEMG data of size (800, 4, N), where 4 is the number of EMG channels, 800 is the timesteps, and N is the number of training datasets. These datasets were stacked and fed into the GRU model for training. The GRU model’s output includes joint kinematics and external load, with dimensions (800, 7, N), representing elbow and shoulder joint angles, velocities, accelerations, and external load. These predicted values are then used as inputs for the inverse dynamics equations of motion to predict joint torques, resulting in dimensions (800, 2, N) for elbow and shoulder torques. In a single training loop, the GRU iterates through each time step of an sEMG input sequence. It computes update and reset gates, calculates the candidate hidden state, and combines them to determine the current hidden state. These hidden states are then used to generate output, such as joint kinematic parameters and external load. Following this, the loss is computed by comparing the model’s output with the actual targets. Through backpropagation through time, gradients of the loss with respect to GRU parameters are calculated, considering time step dependencies. These gradients update the GRU parameters using the Adam optimization algorithm. This loop repeats over multiple epochs and batches to train the GRU model effectively.
3.3 Hyper Parameters of Model
In our model, hyperparameter tuning was crucial. We used a linear/grid search approach to optimize parameters. The model comprises two stacked GRU cells processing EMG signals. We found that using 64 neurons in each hidden layer yielded optimal results. Adam optimization with a learning rate of 0.0001 was chosen for faster convergence. To prevent overfitting, we applied a dropout probability of 0.2 and introduced a physics loss weighting factor of 0.001, determined via linear search. We set the number of epochs to 2000, noting convergence after 1500 epochs due to Adam optimizer’s efficiency.
3.4 Joint torque prediction and evaluation criteria
In this proposed framework, our main focus is to predict joint torque. After completing the training of the PiGRN model, the joint torque computation block is connected with the trained PiGRN block (see Fig. 4). To predict joint torque, testing sEMG data is passed through the trained model to predict joint kinematic parameters and external mass. These predicted parameters are then used as inputs to the inverse dynamic equation of motion to predict joint torques. To evaluate the performance of the model, Percentage root mean square error (RMSE) and the Pearson coefficient of correlation are used, as described in Equation 8 and Equation 10, respectively. % rmse has been used to assess the relative quality of prediction based on its RMSE score.
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (8)
%$\%RMSE=\frac{RMSE*100}{max(y_{i})}$ (9)
Where $n$ is the number of observations or data points, $y_{i}$ represents the actual value (ground truth) for the $i$-th observation, $\hat{y}_{i}$ is the predicted value for the $i$-th observation, and $\sum$ is the summation symbol, indicating a sum over all $n$ observations.
$r=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}}$ (10)
Where $n$ is the number of paired observations, $x_{i}$ represents the value of the $i$-th observation for variable $x$, $\bar{x}$ is the mean (average) value of $x$, $y_{i}$ is the value of the $i$-th observation for variable $y$, $\bar{y}$ is the mean (average) value of $y$, and $\sum$ is the summation symbol, indicating sum over all $n$ observations.
4 RESULTS
In this section, we evaluate the performance of the proposed framework for elbow and shoulder joints by comparing it with selected baseline methods. We begin by describing the training process of the proposed framework.
Figure 5: Loss curve during training
Subsequently, we conduct comparisons to showcase the predicted results of the proposed framework and the baseline methods, including representative outcomes for predicted joint torques, joint kinematic parameters, and external load, as well as detailed and average predictions for five healthy subjects. Finally, we consider the long sequence scenario to assess the robustness and generalization performance of the proposed framework.
(a) 0KG load
(b) 2KG load
(c) 4KG load
(d) 0KG load
(e) 2KG load
(f) 4KG load
Figure 6: Plots for Elbow joint angle and Elbow joint torque prediction of a subject at different loading at hand. Units of elbow torque is N-m where as elbow angles are in normalized in range (0-1) whose maximum value is about ${130}^{0}$.Solid line represents mean value of predicted &\&& true value of all testing dataset whereas shaded region represents spread of them.
Subject 1Subject 2Subject 3Subject 4Subject 5Mean444466668888101010101212121214141414161616164.564.564.564.564.214.214.214.213.53.53.53.54.014.014.014.014.64.64.64.64.024.024.024.027.987.987.987.984.564.564.564.565.65.65.65.6101010107.577.577.577.577.227.227.227.229.539.539.539.535.365.365.365.3615.615.615.615.612.4412.4412.4412.4414.4314.4314.4314.4311.411.411.411.48.268.268.268.267.267.267.267.2614.5614.5614.5614.5612.4412.4412.4412.4415.615.615.615.611.411.411.411.4RMSE in percentageElbow Angle Elbow Torque Shoulder Angle Shoulder Torque
Figure 7: RMSE scores for prediction of joint angle &\&& torque for five different subjects using PiGRN model.
(a) (1.0, 0.21)
(b) (0.1, 0.17)
(c) (0.01, 0.13)
(d) (0.05, 0.065)
(e) (0.001, 0.046)
(f) (0.0001, 0.058)
(g) (1.0, 0.16)
(h) (0.1, 0.15)
(i) (0.01, 0.14)
(j) (0.005, 0.096)
(k) (0.001, 0.075)
(l) (0.0001, 0.099)
Figure 8: Representative comparison plots for elbow angle and torque for different values of physics loss weighting factor ($\lambda_{\text{physics}}$). Tuple in caption represents ($\lambda_{\text{physics}}$, RMSE score)
(a) 0KG load
(b) 2KG load
(c) 4KG load
(d) 0KG load
(e) 2KG load
(f) 4KG load
Figure 9: Plots for Shoulder joint angle and Shoulder joint torque prediction of a subject at different loading at hand. Units of shoulder torque is N-m where as shoulder angles are normalized in range of (0-1) whose maximum value is about ${32}^{0}$.Solid line represents mean value of predicted &\&& true value of all testing dataset whereas shaded region represents spread of them.
Subject 1Subject 2Subject 3Subject 4Subject 5Mean0.40.40.40.40.60.60.60.60.80.80.80.811110.970.970.970.970.990.990.990.990.980.980.980.980.970.970.970.970.990.990.990.990.980.980.980.980.950.950.950.950.990.990.990.990.990.990.990.990.880.880.880.880.980.980.980.980.960.960.960.960.940.940.940.940.790.790.790.790.710.710.710.710.970.970.970.970.930.930.930.930.870.870.870.870.960.960.960.960.960.960.960.960.840.840.840.840.960.960.960.960.910.910.910.910.920.920.920.92Coefficient of CorrelationElbow AngleElbow TorqueShoulder Angle Shoulder Torque
Figure 10: Coefficient of correlation scores for prediction of Elbow &\&& Shoulder joint angles &\&& joint torques for five different subjects using PiGRN model.
(a) Elbow joint angle
(b) Elbow joint angular velocity
(c) Elbow joint angular acceleration
(d) Elbow joint angular acceleration
Figure 11: Elbow Joint angle, angular velocity, angular acceleration &\&& torque prediction for long sequence data
4.1 Evaluation of PiGRN training process
To illustrate the convergence of the proposed PiGRN framework, we present the convergence of the combined total loss, which encompasses both data loss and physics loss for the elbow, shoulder joint, and external load during training, as illustrated in Fig. 5. When predicting the joint kinematic parameters for the elbow, shoulder joint, and external load, the total loss significantly decreases after 150 epochs, displaying minor local oscillations as training progresses, and fully converges after 1250 epochs. We attribute this behavior primarily to the batch size being set to 1 during GRU training. This batch size improves the GRU’s ability to learn the temporal feature of sEMG data more effectively, while causing some oscillations.
4.1.1 Overall comparison
The effectiveness of the proposed method in predicting joint angles and torque using the testing dataset was assessed to verify its robustness. Figures 6 and 9 showcase the outcomes of the proposed framework for elbow and shoulder joints, displaying both joint torques and angles.
ANNRNNGRUPiGRN55551010101015151515202020209.99.99.99.911.611.611.611.69.759.759.759.754.024.024.024.0211.711.711.711.79.49.49.49.49.49.49.49.47.227.227.227.2214.214.214.214.214.614.614.614.612.412.412.412.411.411.411.411.411.611.611.611.612.712.712.712.712.712.712.712.711.411.411.411.4RMSE in percentageElbow AngleElbow TorqueShoulder AngleShoulder Torque
(a) Performance of other models in prediction of joint angle & torques and their comparison with the current model (PiGRN)
0 KG2 KG4 KG444466668888101010104.564.564.564.566.566.566.566.565.165.165.165.167.987.987.987.988.378.378.378.377.457.457.457.459.539.539.539.538.258.258.258.257.627.627.627.628.238.238.238.239.19.19.19.18.78.78.78.7RMSE in percentageElbow AngleElbow TorqueShoulder AngleShoulder Torque
(b) RMSE scores for prediction of joint angle & torque of a subject across different loadings at hand
Figure 12: Comparison of RMSE scores between different models and different loadings
As evidenced in these figures, the predicted values closely match the actual ones, underscoring the remarkable tracking accuracy of the proposed approach. Furthermore, Figure 7 presents the RMSE values (in %) across five subjects, while Figure 10 illustrates the Pearson coefficient of correlation, comparing the predicted elbow angle, elbow torque, shoulder angle, and shoulder torque of the proposed framework with those of baseline methods across different subjects. These figures also presents the mean RMSE values for all five subjects.
4.1.2 Comparison across different load
To investigate the performance of the proposed framework across different loading scenarios, we evaluate the proposed framework by constructing a testing dataset consists of 0, 2, 4kg load sEMG data. Figure 12(b) demonstrates the detailed result of % RMSE scores for one subject, we can find that the proposed framework can achieve satisfactory performance.
4.1.3 Comparison with different methods
To evaluate the effectiveness of the proposed framework quantitatively, we conduct comprehensive comparisons with baseline methods, as depicted in Figure 12(a). The findings reveal that our framework consistently yields lower % RMSEs when forecasting elbow angle, elbow torque, shoulder angle, and shoulder torque, indicating its robustness.
Notably, when compared to deep learning techniques like GRU, RNN, and ANN, our framework demonstrates superior performance owing to its innate ability to automatically extract sophisticated features from the input data. Among these deep learning approaches, PiGRN emerges as the top performer, which we attribute to the integration of fundamental physics principles that regulate the GRU mechanism in our framework. This incorporation enhances performance beyond traditional MSE loss.
4.2 Evaluation of external load and joint kinematic parameters prediction
Figure 11 illustrate the representative results of our proposed framework in predicting joint kinematic parameters. These figures demonstrate the capability of our framework in extracting features related to muscle loading, as well as joint angular velocities and accelerations, from diverse load sEMG data.
4.3 Evaluation of long sequence joint angle and torque prediction
To assess the flexibility of our proposed framework in handling varying input sequence lengths and corresponding output sequences, Figure 11 demonstrates long sequence predictions of elbow joint angle and joint torque. These results indicate that our trained model is capable of processing input sequences of different lengths beyond those it was initially trained on.
4.4 Effect of physics informed weighting factor ($\lambda_{\text{physics}}$)
We conducted an investigation into the impact of the physics loss weighting factor ($\lambda_{\text{physics}}$) on the performance of our proposed framework, focusing on scenarios involving elbow joint angle and torque. The detailed results are presented in Figure 8(l). Specifically, we examined six physics-informed weighting factors: 1.0, 0.1, 0.01, 0.05, 0.001, and 0.0001, while keeping other hyperparameters constant. Analysis of Figure 8(l) reveals that the proposed framework achieves optimal performance with a smaller weighting factor of 0.001.
5 DISCUSSION AND FUTURE WORK
In this research, we employed a sophisticated physics-informed gated recurrent network (PiGRN), combining a GRU time-series architecture with the differential equation of motion for the upper limb. This hybrid model predicts joint kinematic parameters at the elbow and shoulder joints and the external load, which are then used to predict joint torque. Unlike ANN or CNN models, our framework allows flexible input and output, making it suitable for real-time joint torque prediction in HMI development. During training, we used 800 timesteps of data to speed up the process. However, once trained, the model can handle input sequences of any length. Figure 11 shows the prediction of joint elbow angle and torque for 5000 timesteps, demonstrating the PiGRN model’s generalization capability. Currently, we predict the elbow and shoulder joints with two degrees of freedom (DOF). Future work will expand to more joints and DOF, potentially covering the entire upper or lower limb.
For the physics-informed component, we utilized the equations of motion, where the joint angles, angular velocities, and angular accelerations serve as inputs. These inputs are predicted from the GRU component itself. This approach reduces the time needed for automatic differentiation as well. Figure 11 showcases a prediction plot for elbow angle, angular velocity, and elbow angular acceleration as an illustrative example. From the figure, it is evident that the predictions closely match the actual values, exhibiting sharp patterns. Thus, the proposed PiGRN model successfully captures the intricate patterns of joint kinematic parameters once trained, incorporating physical laws into its architecture and training process. This ensures that the network’s predictions align with established physical principles. It is worth noting that in this study, actual values for the joint kinematic parameters were computed using the central difference method. However, in future applications, joint angular accelerations can be measured experimentally using sensors such as inertial measurement units (IMUs), from which angular velocities can be computed.
During our framework implementation, we found that predicting shoulder angle and torque isn’t as accurate as elbow angle and torque, resulting in higher RMSE scores (Fig. 7). Our focus was on predicting multi-joint dynamics in sagittal plane movement only, where shoulder angle was smaller than elbow angle during the biceps curl-like motion. For this, we used sEMG from the biceps and triceps, neglecting deltoid muscle involvement. Future work could include deltoid muscle sEMG to improve shoulder prediction accuracy.
6 CONCLUSION
This paper presents a novel physics-informed gated recurrent neural network (PiGRN) developed to predict joint torque using sEMG data under various load conditions. The proposed method is capable of predicting multi-joint kinematic parameters and external loads from sEMG sequences of different lengths. Unlike conventional deep learning approaches, our framework uniquely integrates predicted joint kinematic parameters and external loads into physics-based domain knowledge as soft constraints to regularize the GRU’s loss function, a technique that has not been implemented before. This strategy reduces the dependence on automatic differentiation and decreases training time. Experiments with sEMG data across three different load groups demonstrate the method’s effectiveness in predicting joint torque and kinematics. We believe this framework could serve as a general method for predicting joint torque and kinematics, making it well-suited for real-time applications in the development of robust HMIs. Although this study focused on data involving two degrees of freedom, two joints, and in-plane movements, future work will aim to address more generalized datasets and musculoskeletal models, as well as explore unsupervised learning techniques.
Acknowledgment
The authors would like to thank Anant Jain from Indian Institute of Technology (IIT) Delhi for helping in data collection. This study is partly supported by the Joint Advanced Technology Centre (JATC) (Grant no.: RP03830G) and the I-Hub Foundation for cobotics (IHFC) section-8 company (Grant no.: GP/2021/RR/010) at IIT Delhi, sponsored by the Ministry of Education (MoE), Govt. of India.